an-3an-1-4an-2=4.3n


1
Expert's answer
2021-07-22T06:01:04-0400
"a_n-3a_{n-1}-4a_{n-2}=4\cdot3^n"

The associated linear homogeneous equation is 


"a_n-3a_{n-1}-4a_{n-2}=0"

The characteristic equation is


"r^2-3r-4=0"

"(r+1)(r-4)=0"

"r_1=-1, r_2=4"

The solution is


"a_n^{(h)}=\alpha_1(-1)^n+\alpha_2(4)^n"



Because "F(n)=4\cdot3^n," a reasonable trial solution is


"a_n^{(p)}=C\cdot3^n,"

where "C" is a constant.

Substitute


"C\cdot3^n-3C\cdot3^{n-1}-4C\cdot3^{n-2}=4\cdot3^n"

"9C-9C-4C=36"

"C=-9"

Hence the particulr solution is


"a_n^{(p)}=-9\cdot3^n"

All solutions are of the form


"a_n=\alpha_1(-1)^n+\alpha_2(4)^n-9\cdot3^n"

where "\alpha_1" and "\alpha_2" are a constants.